Overhead Variances, Two- And Three-Variance Analyses

Oerstman, Inc., uses a standard costing system and develops its overhead rates from the current annual budget. The budget is based on an expected annual output of 123,000 units requiring 492,000 direct labor hours. (Practical capacity is 512,000 hours.) Annual budgeted overhead costs total $752,760, of which $546,120 is fixed overhead. A total of 119,000 units using 490,000 direct labor hours were produced during the year. Actual variable overhead costs for the year were $240,000, and actual fixed overhead costs were $555,450.

Required:

1. Compute overhead variances using a two-variance analysis.

Budget Variance $ Unfavorable
Volume Variance $ Unfavorable

2. Compute overhead variances using a three-variance analysis.

Spending Variance $ Unfavorable
Efficiency Variance $ Unfavorable
Volume Variance $ Unfavorable
Answer #1

1. A) Budget Variance

Budget variance= actual overhead- budgeted overhead

Budgeted overhead= fixed overhead + budgeted variable overhead for actual production

Fixed overhead= 546120

Budgeted variable overhead for actual production= total units X standard hours per unit X Standard rate of variable overhead

Standard hours per unit= budgeted direct labour hours / expected annual units = 492000/123000= 4

Standard rate of variable overhead =annual budgeted overhead – fixed overhead /budgeted direct labour hours = 752760 – 546120 / 492000 = 0.42

standard rate of fixed overhead = fixed overhead / budgeted direct labour hours = 546120/492000= 1.11

Budgeted overhead= 546120 + ( 119000 X 4 X 0.42) = 746040

Actual overhead = actual variable overhead + actual fixed overhead= 240000+ 555450 = 795450

Budgeted variance = 795450- 746040 = 49410 U

b) volume variance = fixed overhead applied - budgeted fixed overhead

fixed overhead applied = total units X Standard hours per unit X standard rate of fixed overhead = 119000 X 4 X 1.11 = 528360

volume variance = 528360 – 546120 = 17760 U

2. A) Spending variance = variable overhead rate variance + fixed overhead budget variance

variable overhead rate variance = Standard rate – actual rate

actual rate of variable overhead= variable overhead / actual direct labour hours = 240000/490000 =0.49 per hour

variable overhead rate variance = (Standard rate of variable overhead – Actual rate of variable overhead ) X actual direct labour hours= 0.42 – 0.49=0.07 X 490000 = 34300 U

fixed overhead budget variance = budgeted fixed overhead – actual fixed overhead = 546120 – 555450 = 9330 U

Spending variance = 34300 U + 9330 U= 43630

B) Variable overhead efficiency variance = Standard Hours – actual Hours X standard rate = (119000 X 4) – 490000 X 0.42 = 5880 U

C) volume variance = fixed overhead applied – budgeted fixed overhead = 528360 – 546120 = 17760 U

Note : ‘U’ indicated the Unfavorable